find w(t) = p(x,y) on a unit circle that corresponds to t=5*pi/6
thanks(:

Answer by stanbon(75887)

find w(t) = p(x,y) on a unit circle that corresponds to t=5*pi/6
---
Well, I guess your t is an angle of (5/6)pi = 150 degrees.
"That was not clear in your first posting".
---------------------------------
That point on the unit circle would be (x,y)
where x = cos[(5/6)pi) = -(1/2)sqrt(3)
and y = sin[(5/6)pi] = 1/2
=======
========================================
Cheers,
Stan H.